#2, ch 7

Required information [The following information applies to the questions displayed below.] The following information relates to a joint production process for three products, with a total joint production cost of $135,000. There are no separable processing costs for any of the three products. Product Sales Value at Split-Off Units at Split-Off 1 $ 189,000 260 2 54,000 780 3 27,000 1,560 $ 270,000 2,600 What percentage of joint cost is allocated to each of the three products using the physical units method?
Answer #1
Product Allocated joint cost
1 135,000*260/2600 = 13,500
2 135,000*780/2600= 40,500
2 135,000*1560/2600=81,000

Comment if you face any issues